Circular No. 5725

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET SHOEMAKER-LEVY (1993e) Cometary images have been discovered by C. S. Shoemaker, E. M.

Shoemaker and D. H. Levy on films obtained with the 0.46-m Schmidt telescope at Palomar. The appearance was most unusual in that the comet appeared as a dense, linear bar ∼ 1′ long and oriented roughly east-west; no central condensation was observable, but a fainter, wispy ‘tail’ extended north of the bar and to the west. The object was confirmed two nights later in Spacewatch CCD scans by J. V. Scotti, who described the nuclear region as a long, narrow train ∼ 47′′ in length and ∼ 11′′ in width, aligned along p.a. 80^{o}-260^{o}. At least five discernible condensations were visible within the train, the brightest being ∼ 14′′ from the southwestern end. Dust trails extended 4′.20 in p.a. 74^{o} and 6′.89 in p.a. 260^{o}, roughly aligned with the ends of the train and measured from the midpoint of the train. Tails extended > 1′ from the nuclear train, the brightest component extending from the brightest condensation to 1′.34 in p.a. 286^{o}. The measurements below refer to the midpoint of the bar or train.

1993 UT  α_{2000}  δ_{2000}  m_{1} Observer
Mar. 24.35503 12^{h}26^{m}39.^{s}27 −4^{o}03′32′′.9 14  Shoemaker
24.43072 12 26 37.21 −4 03 23.0  _{”}
26.29531 12 25 42.24 −3 57 55.7 13.9 Scotti
26.30479 12 25 42.09 −3 57 53.7 16.7  _{”}
26.31448 12 25 41.63 −3 57 53.7  _{”}
26.41291 12 25 38.70 −3 57 34.8  _{”}

C. S. Shoemaker, E. M. Shoemaker, D. H. Levy and P. Bendjoya (Palomar).

Measurers D. H. Levy, J. Mueller, P. Bendjoya and E. M. Shoemaker. J. V. Scotti (Kitt Peak). Last observation made through cirrus.

The comet is located ∼ 4^{o} from Jupiter, and the motion suggests that it may be near Jupiter’s distance.

SUPERNOVA 1993E IN KUG 0940+495 D. D. Balam and G. C. L. Aikman report a measurement of V =

20.3 ± 0.1 and B − V = +0.51 on Feb. 26.28 UT, using the 1.85-m reflector (+ CCD) at the Dominion Astrophysical Observatory.

1993 March 26 Brian G. Marsden